# FREA: Feasibility-Guided Generation of Safety-Critical Scenarios with Reasonable Adversariality

### v1

@@ Abstract
Generating safety-critical scenarios, which are essential yet difficult to collect at scale, offers an effective method to evaluate the robustness of autonomous vehicles (AVs). Existing methods focus on optimizing adversariality while preserving the naturalness of scenarios, aiming to achieve a balance through data-driven approaches. However, without an appropriate upper bound for adversariality, the scenarios might exhibit excessive adversariality, potentially leading to unavoidable collisions. In this paper, we introduce FREA, a novel safety-critical scenarios generation method that incorporates the Largest Feasible Region (LFR) of AV as guidance to ensure the REasonableness of the Adversarial scenarios. Concretely, FREA initially pre-calculates the LFR of AV from offline datasets. Subsequently, it learns a reasonable adversarial policy that controls critical background vehicles (CBVs) in the scene to generate adversarial yet AV-feasible scenarios by maximizing a novel feasibility-dependent objective function. Extensive experiments illustrate that FREA can effectively generate safety-critical scenarios, yielding considerable near-miss events while ensuring AV’s feasibility. Generalization analysis also confirms the robustness of FREA in AV testing across various surrogate AV methods and traffic environments.
@@ 1 Introduction
Rapid advancements in AI-driven autonomous driving technologies have enabled autonomous vehicles (AVs) to operate efficiently, safely, and comfortably in most daily scenarios [1]. However, their performance in the complex real-world remains uncertain due to the existence of safety-critical yet rare scenarios, which are hard to collect at scale. Accurately generating these safety-critical scenarios and evaluating the AV performance in these scenarios is crucial for advancing the deployment of autonomous driving systems [2].
@@
Inspired by the concept of feasibility in Safe Reinforcement Learning (Safe RL) [13, 14, 15], we propose an innovative approach using the feasibility of the AV to define the upper bound for adversariality in safety-critical scenarios. Specifically, under a given critical scenario, if an optimal policy exists that allows the AV to avoid potential collisions, then the current state of the AV is within its LFR, as detailed in Section 3.1. Therefore, we naturally use the boundary of the AV’s LFR as the upper bound for the scenario’s adversariality to ensure that even the most adversarial scenario generated still satisfies the AV’s feasibility.
@@
In this paper, we propose FREA, a novel method designed to generate adversarial yet AV-feasible scenarios. To address the “curse of dimensionality”[16], similar to approaches in [3, 4], we manipulate the behavior of some critical background vehicles (CBVs) at pivotal moments to reduce computational complexity. Moreover, we incorporate the AV’s feasibility as guidance to ensure that the adversariality of CBV policy remains within a reasonable range. Specifically, FREA involves a two-stage framework: initially, we pre-train the optimal feasible value network of AV using offline datasets; subsequently, we introduce a feasibility-dependent objective that enables CBV to learn a reasonable adversarial policy. Experimental results demonstrate that feasibility-guided adversarial policy facilitates the generation of near-miss events while significantly reducing over-adversarial scenarios and unavoidable collisions. Further performance evaluation of AVs confirms the robustness and generalization ability of FREA in AV testing across various surrogate AV methods and environments. The main contributions of this work are concluded as follows:
@@
    We propose FREA, a safety-critical scenario generation method that leverages the Largest Feasible Region (LFR) of AV as guidance to ensure the reasonableness of the adversarial scenarios.
@@
    We introduce a feasibility-dependent objective that enables CBVs to learn a reasonable adversarial policy, thereby generating adversarial yet AV-feasible, safety-critical scenarios.
  3. 3.
@@ 2 Related Work
Safety-Critical Scenarios Generation. Recently, the Deep Generative Model (DGM) method has gained significant attention in scenario generation for its ability to construct neural networks that model data distributions and generate scenarios by sampling from these models [6, 7, 9, 10, 17]. Commonly, these methods output vehicle trajectories that may not always adhere to dynamic constraints or exhibit the necessary smoothness, raising concerns about their dynamic authenticity. In contrast, optimization-based methods [5, 11, 18] optimize the actions of all vehicles in a scene, ensuring that these actions stay within predefined constraints and maintain continuity to guarantee authenticity. These methods typically incorporate various regularization terms to enhance the rationality of the actions. However, this approach often leads to the “curse of dimensionality” [16] by simultaneously optimizing actions for all vehicles. RL-based strategies [3, 4, 12] address this by controlling only critical vehicles at crucial moments, aiming to enhance adversariality while ensuring consistency with natural dataset distributions. These methods typically evaluate adversariality using metrics like collision rates or the minimum distances to the AV. Although they emphasize the naturalness of macro-scenarios, the pursuit of maximum adversariality still can result in excessive adversarial scenarios. To mitigate these concerns, we introduce a feasibility-guided adversarial policy of CBV that ensures AV’s feasibility while optimizing scenario adversariality.
@@
Safe RL with Hard Constraint. Safe RL is usually modeled as a Constrained Markov Decision Process (CMDP) defined by $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,h,c,\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ represent the state and action spaces, $P$ describes the transition dynamics, $r$ is the reward function, $h$ measures constraint violations, $c$ is a bounded cost function $[0,C_{\max}]$, and $\gamma$ is the discount factor within $(0,1)$. The goal is to optimize a policy $\pi$ to maximize rewards while adhering to safety constraints, especially zero state-wise constraint violations $(h(s_{t})\leq 0,a\sim\pi,\forall t\in\mathbb{N})$, critical in domains like autonomous driving. To enhance policy safety, various strategies have been developed: [19] proposes a safety index that ensures persistent safety for energy-dissipating policies, while [20, 13, 14] use reachability certificates to manage the worst-case constraint value and constrain it below zero in online policy optimization. [15] decouples the learning of the reachability certificate from policy optimization, conducting it offline to achieve a more stable training process. Since we require the LFR of the AV and the online training with fixed-policy AV only yields a subset of LFR, following [15], we adopt an offline method to learn the AV’s LFR, which also provides a much more stable training process.
@@ 3 Methodology
In this section, we explore the integration of the AV’s feasibility into adversarial policy learning for the CBV. Initially, the AV’s optimal feasible value network is trained offline (Section 3.1) to determine the LFR of the AV. Subsequently, under the proper assumption, CBV can access the AV’s LFR via its state. Building on this, we propose a feasibility-dependent adversarial objective for the CBV, designed to create scenarios with reasonable adversariality (Section 3.2).
To clarify the notation in CMDP, in the remainder of this paper, the elements associated with AV are added with superscripts $\cdot^{\text{AV}}$, while those associated with CBV have no special superscripts.
@@ 3.1 Largest Feasible Region of AV
Determining the LFR of the AV is crucial for the success of feasibility-guided adversarial policy learning for the CBV. Given the efficacy of Hamilton-Jacobi Reachability (HJR) [21] in the domain of Safe RL, we adopt the HJR framework to instantiate LFR. First, we provide a brief overview of the basic concepts of HJR (Definition 1) and then dive into the detailed instantiation of the LFR under the HJR framework (Definition 2).
@@
Based on [21], the optimal feasible state-value function $V^{*}_{h}$ and the optimal feasible action-value function $Q^{*}_{h}$ are defined in Eqs. 1 and 2.
@@
  • $V_{h}^{*}(s^{\text{AV}})\leq 0\Rightarrow\exists\pi^{\text{AV}},\max_{t\in\mathbb{N}}h\left(s_{t}^{\text{AV}}\right)\leq 0$, that is, when $V_{h}^{*}(s^{\text{AV}})$ is not larger than zero, there exists an optimal policy for the AV that satisfies the hard constraints.
  • $V_{h}^{*}(s^{\text{AV}})>0\Rightarrow$ $\forall{\pi^{\text{AV}}},\max_{t\in\mathbb{N}}h\left(s_{t}^{\text{AV}}\right)>0$, that is, when $V_{h}^{*}(s^{\text{AV}})$ is larger than zero, no policy can ensure the safety of AV under the hard constraints.
Therefore, the optimal feasible value function serves as an indicator of whether there exists a policy of AV that achieves hard constraints. Consequently, we can easily derive the following definition of the Largest Feasible Region (LFR).
@@
To accurately obtain the LFR, we adopt the method proposed in [15], which involves pre-collecting driving trajectories using different AV algorithms and then deriving the LFR through offline learning. The optimal feasible value functions $V_{h}^{*}(s^{\text{AV}})$ and $Q_{h}^{*}(s^{\text{AV}})$ can be obtained, respectively, by minimizing the following objective functions:
$\mathcal{L}_{V_{h}}$ $=\mathbb{E}_{\mathcal{D}}\left[L_{\text{rev}}^{\tau}\left(Q_{h}(s^{\text{AV}},a^{\text{AV}})-V_{h}(s^{\text{AV}})\right)\right],$ (3)
$\mathcal{L}_{Q_{h}}$ $=\mathbb{E}_{\mathcal{D}}\left[\left(\left((1-\gamma)h(s^{\text{AV}})+\gamma\max\left\{h(s^{\text{AV}}),V_{h}\left({s^{\text{AV}}}^{\prime}\right)\right\}\right)-Q_{h}(s^{\text{AV}},a^{\text{AV}})\right)^{2}\right],$ (4)
where $L_{\mathrm{rev}}^{\tau}(u)=|\tau-\mathbb{I}(u>0)|u^{2}$ for $\tau\in(0.5,1)$ and $\mathcal{D}:=\left(s^{\text{AV}},a^{\text{AV}},{s^{\text{AV}}}^{\prime}\right)$ is the offline dataset, comprising both safe and unsafe trajectories. Further dataset collection details are in Section C.1.
With the well-trained $V_{h}^{*}(s^{\text{AV}})$, we can easily assess whether the AV is within the LFR based on its current state $s^{\text{AV}}$. Additionally, by interpreting the AV’s state as a scenario description and applying the viewpoint transformation function $g(\cdot)$, we derive $s^{\text{AV}}$ from $g(s)$, where $s$ is the CBV’s state. This links the AV’s LFR to the CBV’s adversarial policy learning, showing $V^{\ast}_{h}(s^{\text{AV}})=V^{\ast}_{h}(g(s))$, thus connecting AV’s feasibility with CBV strategy.
@@ 3.2 Feasibility-Dependent Adversarial Objective of CBV
With the well-trained LFR of AV from Section 3.1 and the relation between $s^{\text{AV}}$ and $s$, the CBV can access AV’s LFR through $s$. This enables us to form a standard CMDP framework for the CBV.
The standard feasibility-dependent objective function usually adopts distinct objects in the feasible and infeasible regions, with the principle of maximizing expected cumulative reward in the feasible region while minimizing constraint violations as much as possible in the infeasible region [13, 14, 15]. With the predetermined LFR and the state-action pair from the rollout buffer, we modify the on-policy version of the feasibility-dependent objective function from [13], incorporating a stricter constraint, as shown in Eqs. 5 and 6 (Section A.1 demonstrates the theoretical relationships between Eq. 6 and the original advantage function in [13]). Intuitively, since the advantage describes the superiority of taking action $a$ at the current state $s$, it inspires us to impose stricter restrictions on the region where reward advantage applies, considering not only the current state’s feasibility but also the feasibility of the next state.
$L(\theta)=$ ${\mathbb{E}}_{\pi_{\theta_{k}}}\left[\min\left(r_{t}(\theta)\overline{A^{\pi_{\theta_{k}}}}(s,a),\operatorname{clip}\left(r_{t}(\theta),1-\epsilon,1+\epsilon\right)\overline{A^{\pi_{\theta_{k}}}}(s,a)\right)\right],$ (5)
$\overline{A^{\pi_{\theta_{k}}}}(s,a)=$ $A_{r}^{\pi_{\theta_{k}}}(s,a)\cdot\mathbb{I}_{V^{\ast}_{h}(g(s))\leq 0,\,V^{\ast}_{h}(g(s^{\prime}))\leq 0}+A_{h}^{*}(s,a)\cdot(1-\mathbb{I}_{V^{\ast}_{h}(g(s))\leq 0,\,V_{h}^{\ast}(g(s^{\prime}))\leq 0}),$ (6)
where $r_{t}(\theta)=\frac{\pi_{\theta}(a\mid s)}{\pi_{\theta_{k}}(a\mid s)}$, $A_{r}^{\pi_{\theta_{k}}}(s,a)$ is CBV’s adversarial reward advantage, and $A_{h}^{*}\left(s,a\right)$ is AV’s feasibility advantage function, but accessed through CBV’s state and action.
Since $Q_{h}^{*}\left(s^{\text{AV}},a^{\text{AV}}\right)$ is related to $a^{\text{AV}}$ instead of $a$, we cannot directly derive $A^{*}_{h}\left(s,a\right)$ in the same form as $A_{r}^{\pi_{\theta_{k}}}(s,a)=Q_{r}^{\pi_{\theta_{k}}}(s,a)-V_{r}^{\pi_{\theta_{k}}}(s)$. Fortunately, the constraint function $h(\cdot)$ is solely related to the state; under the appropriate assumption, we can derive the following lemma:
@@
As the BVs follow deterministic policy, the optimal feasible action-value function of AV can be achieved by CBV’s current state and next state (see Section A.2 for proof).
$Q_{h}^{*}\left(s^{\textnormal{AV}},a^{\textnormal{AV}}\right)=Q_{h}^{*}\left(g\left(s\right),g\left(s^{\prime}\right)\right)=\begin{cases}V_{h}^{*}\left(g\left(s^{\prime}\right)\right)&h\left(g\left(s^{\prime}\right)\right)\geq h\left(g\left(s\right)\right)\\ \max\{h\left(g\left(s\right)\right),V_{h}^{*}\left(g\left(s^{\prime}\right)\right)\}&h\left(g\left(s^{\prime}\right)\right)<h\left(g\left(s\right)\right)\end{cases}$ (7)
With Lemma 1 and $V_{h}^{*}\left(s^{\text{AV}}\right)=V_{h}^{*}\left(g(s)\right)$, we can derive AV’s feasibility advantage through CBV’s trajectory as:
$A_{h}^{*}\left(g\left(s\right),g\left(s^{\prime}\right)\right)=\begin{cases}V_{h}^{*}\left(g\left(s^{\prime}\right)\right)-V_{h}^{*}\left(g\left(s\right)\right)&h\left(g\left(s^{\prime}\right)\right)\geq h\left(g\left(s\right)\right)\\ \max\{h\left(g\left(s\right)\right),V_{h}^{*}\left(g\left(s^{\prime}\right)\right)\}-V_{h}^{*}\left(g\left(s\right)\right)&h\left(g\left(s^{\prime}\right)\right)<h\left(g\left(s\right)\right)\end{cases}$ (8)
By addressing all the missing parts in Eq. 6, we can replace the original advantage in PPO [22] and formulate the overall algorithm of feasibility-guided adversarial policy, as shown in Algorithm 1.
@@
8:Online Part
9:Initialize policy parameters $\theta_{0}$, reward value function parameters $\omega_{0}$
@@
11:     Collect set of trajectories $\mathcal{D}_{k}=\{\tau_{i}\}$ with policy $\pi_{\theta_{k}}$, where $\tau_{i}$ is a $T$-step episode.

### v2

@@ Abstract
Generating safety-critical scenarios, which are essential yet difficult to collect at scale, offers an effective method to evaluate the robustness of autonomous vehicles (AVs). Existing methods focus on optimizing adversariality while preserving the naturalness of scenarios, aiming to achieve a balance through data-driven approaches. However, without an appropriate upper bound for adversariality, the scenarios might exhibit excessive adversariality, potentially leading to unavoidable collisions. In this paper, we introduce FREA, a novel safety-critical scenarios generation method that incorporates the Largest Feasible Region (LFR) of AV as guidance to ensure the REasonableness of the Adversarial scenarios. Concretely, FREA initially pre-calculates the LFR of AV from offline datasets. Subsequently, it learns a reasonable adversarial policy that controls the scene’s critical background vehicles (CBVs) to generate adversarial yet AV-feasible scenarios by maximizing a novel feasibility-dependent adversarial objective function. Extensive experiments illustrate that FREA can effectively generate safety-critical scenarios, yielding considerable near-miss events while ensuring AV’s feasibility. Generalization analysis also confirms the robustness of FREA in AV testing across various surrogate AV methods and traffic environments. For more information, visit the project website: https://currychen77.github.io/FREA/
@@ 1 Introduction
Rapid advancements in AI-driven autonomous driving technologies have enabled autonomous vehicles (AVs) to operate efficiently, safely, and comfortably in most daily scenarios [1]. However, their performance in the complex real world remains uncertain due to safety-critical yet rare scenarios, which are hard to collect at scale. Accurately generating these safety-critical scenarios and evaluating the AV performance is crucial for advancing autonomous driving systems [2].
@@
Inspired by the concept of feasibility in Safe Reinforcement Learning (Safe RL) [13, 14, 15], we propose an innovative approach using the feasibility of the AV to define the upper bound for adversariality in safety-critical scenarios. Specifically, under a given critical scenario, if an optimal policy exists that allows the AV to avoid potential collisions, then the current state of the AV is within its Largest Feasible Region (LFR), as detailed in Section 3.1. Therefore, we naturally use the boundary of the AV’s LFR as the upper bound for the scenario’s adversariality to ensure that even the most adversarial scenario generated still satisfies the AV’s feasibility.
@@
In this paper, we propose FREA, a novel method designed to generate adversarial yet AV-feasible scenarios. To address the “curse of dimensionality” [16], similar to approaches in [3, 4], we manipulate the behavior of some critical background vehicles (CBVs) at pivotal moments to reduce computational complexity. Moreover, we incorporate the AV’s feasibility as guidance to ensure that the adversariality of CBV policy remains within a reasonable range. Specifically, FREA involves a two-stage framework: initially, we pre-train the optimal feasible value network of AV using offline datasets; subsequently, we introduce a feasibility-dependent adversarial objective that enables CBV to learn a reasonable adversarial policy. Experimental results demonstrate that feasibility-guided adversarial policy facilitates the generation of near-miss events while significantly reducing over-adversarial scenarios and unavoidable collisions. Further performance evaluation of AVs confirms the robustness and generalization ability of FREA in AV testing across various surrogate AV methods and environments. The main contributions of this work are concluded as follows:
@@
    We propose FREA, a safety-critical scenario generation method that leverages the Largest Feasible Region (LFR) of AV as guidance to ensure the adversarial reasonableness of the Critical Background Vehicles (CBVs), thereby generating adversarial yet AV-feasible scenarios.
@@ 2 Related Work
Safety-Critical Scenarios Generation. Recently, the Deep Generative Model (DGM) method has gained significant attention in scenario generation for its ability to construct neural networks that model data distributions and generate scenarios by sampling from these models [6, 7, 9, 10, 17]. Commonly, these methods output vehicle trajectories that may not always adhere to dynamic constraints or exhibit the necessary smoothness, raising concerns about their dynamic authenticity. In contrast, optimization-based methods [5, 11, 18] optimize the actions of all vehicles in a scene, ensuring that these actions stay within predefined constraints and maintain continuity to guarantee authenticity. These methods typically incorporate various regularization terms to enhance the rationality of the actions. However, with more vehicles, the “curse of dimensionality” [16] often hampers optimization-based methods from finding plausible trajectories, reducing performance [5, 11]. RL-based strategies [3, 4, 12, 19] address this by controlling only Critical Background Vehicles (CBVs) at crucial moments, aiming to enhance adversariality while ensuring consistency with natural dataset distributions. These methods typically evaluate adversariality using metrics like collision rate [3, 4], collision likelihood [19], or the minimum distance [12] to the AV. Although they emphasize macro-scenarios’ naturalness, pursuing maximum adversariality still results in excessive adversarial scenarios. To mitigate these concerns, we introduce a feasibility-guided adversarial policy of CBV that ensures AV’s feasibility while optimizing scenario adversariality.
@@
Safe RL with Hard Constraint. Safe RL is usually modeled as a Constrained Markov Decision Process (CMDP) defined by $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,h,c,\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ represent the state and action spaces, $P$ describes the transition dynamics, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $h:\mathcal{S}\rightarrow\mathbb{R}$ is the constraint violation function, $c:\mathcal{S}\rightarrow\left[0,C_{\max}\right]$ is a bounded cost function, and $\gamma$ is the discount factor within $(0,1)$. Typically, $c(s)=\max\{h(s),0\}$. The goal is to optimize a policy $\pi$ to maximize rewards while adhering to safety constraints, especially worst-case constraints in safety-critical domains like autonomous driving. To enhance safety, various strategies have been developed: [20] proposes a safety index that ensures persistent safety for energy-dissipating policies, while [21, 13, 14] employ reachability certificates to ensure the worst-case constraints remain non-positive during online policy optimization. [15] decouples the learning of the reachability certificate from policy optimization, performing it offline for more stable training. As we need the LFR of the AV and online training with a fixed-policy AV only yields a subset of LFR, following [15], we adopt an offline method to comprehensively learn the AV’s LFR, enhancing training stability.
@@ 3 Methodology
In this section, we integrate AV’s feasibility into adversarial policy learning for the Critical Background Vehicle (CBV), formulating FREA as shown in Algorithm 1. The offline part (Section 3.1) trains the AV’s optimal feasible value network using the offline dataset to approximate the Largest Feasible Region (LFR) of the AV. In the online part (Section 3.2), we propose a feasibility-dependent adversarial objective for CBV training with the predetermined LFR. For clarity, the elements associated with AV are denoted with $\cdot^{\textnormal{AV}}$, while CBV-related elements have no special superscripts.
@@ 3.1 Largest Feasible Region of AV
Given the efficacy of Hamilton-Jacobi Reachability (HJR) [22] in the domain of Safe RL, we adopt the HJR framework to instantiate LFR. First, we provide a brief overview of HJR (Definition 1) and then dive into the detailed instantiation of the LFR under the HJR framework (Definition 2).
@@
Based on [13, 15, 22], the optimal feasible state-value function $V^{*}_{h}$ and the optimal feasible action-value function $Q^{*}_{h}$ are defined in Eqs. 1 and 2.
@@
  • $V_{h}^{*}(s^{\textnormal{AV}})\leq 0\Rightarrow\exists\pi^{\textnormal{AV}},\max_{t\in\mathbb{N}}h\left(s_{t}^{\textnormal{AV}}\right)\leq 0$, that is, when $V_{h}^{*}(s^{\textnormal{AV}})$ is not larger than zero, there exists an optimal policy for the AV that satisfies the hard constraints.
  • $V_{h}^{*}(s^{\textnormal{AV}})>0\Rightarrow$ $\forall{\pi^{\textnormal{AV}}},\max_{t\in\mathbb{N}}h\left(s_{t}^{\textnormal{AV}}\right)>0$, that is, when $V_{h}^{*}(s^{\textnormal{AV}})$ is larger than zero, no policy can ensure the safety of AV under the hard constraints.
Therefore, the optimal feasible value function serves as an indicator of whether there exists a policy of AV that achieves hard constraints. Consequently, we can easily derive the following definition.
@@
To accurately obtain the LFR, we adopt the method proposed in [15], which involves pre-collecting driving trajectories using different AV algorithms and then deriving the LFR through offline learning. The optimal feasible value functions $V_{h}^{*}(s^{\textnormal{AV}})$ and $Q_{h}^{*}(s^{\textnormal{AV}})$ can be obtained, respectively, by minimizing the following objective functions:
$\mathcal{L}_{V_{h}}(\omega)$ $=\mathbb{E}_{\mathcal{D}}\left[L_{\text{rev}}^{\tau}\left(Q_{h}(s^{\textnormal{AV}},a^{\textnormal{AV}};\phi)-V_{h}(s^{\textnormal{AV}};\omega)\right)\right],$ (3)
$\mathcal{L}_{Q_{h}}(\phi)$ $=\mathbb{E}_{\mathcal{D}}\left[\left(\left((1-\gamma)h(s^{\textnormal{AV}})+\gamma\max\left\{h(s^{\textnormal{AV}}),V_{h}({s^{\textnormal{AV}}}^{\prime};\omega)\right\}\right)-Q_{h}(s^{\textnormal{AV}},a^{\textnormal{AV}};\phi)\right)^{2}\right],$ (4)
where $L_{\mathrm{rev}}^{\tau}(u)=|\tau-\mathbb{I}(u>0)|u^{2}$ for $\tau\in(0.5,1)$ and $\mathcal{D}$ is the offline dataset, comprising both safe and unsafe trajectories. Further dataset collection details are in Section C.1.
With the well-trained $V_{h}^{*}(s^{\textnormal{AV}})$, we can easily assess whether the AV is within the LFR based on its current state $s^{\textnormal{AV}}$. Additionally, by interpreting the AV’s state as a scenario description and applying the viewpoint transformation function $g(\cdot)$, we derive $s^{\textnormal{AV}}$ from $g(s)$, where $s$ is the CBV’s state. This links the AV’s LFR to the CBV’s adversarial policy learning, showing $V^{\ast}_{h}(s^{\textnormal{AV}})=V^{\ast}_{h}(g(s))$, thus connecting the AV’s feasibility with CBV policy.
@@ 3.2 Feasibility-Dependent Adversarial Objective of CBV
With the predetermined LFR of AV from Section 3.1 and the relation between $s^{\textnormal{AV}}$ and $s$, the CBV can access AV’s LFR through $s$ to evaluate feasibility. This enables us to form a standard CMDP framework for CBV to encourage adversarial behavior while ensuring reasonableness.
The standard feasibility-dependent objective function usually adopts distinct objects in the feasible and infeasible regions, with the principle of maximizing expected cumulative reward in the feasible region while minimizing constraint violations as much as possible in the infeasible region [13, 14, 15]. With the predetermined LFR and the data $(s,a,s^{\prime},r)$ sampled from the rollout buffer, we modified the on-policy version of the feasibility-dependent objective function from [13], incorporating stricter constraints, as shown in Eqs. 5 and 6. Intuitively, we hope to maximize CBV’s adversarial rewards when the current state $s^{\textnormal{AV}}=g\left(s\right)$ and the next state ${s^{\textnormal{AV}}}^{\prime}=g\left(s^{\prime}\right)$ of the AV are within LFR and minimize the AV’s constraint violations otherwise (Section A.1 demonstrates the theoretical relationships between Eq. 6 and the origin advantage function in [13]).
$L(\theta)=$ ${\mathbb{E}}_{\pi_{\theta_{k}}}\left[\min\left(r_{t}(\theta)A^{\pi_{\theta_{k}}}(s,a),\operatorname{clip}\left(r_{t}(\theta),1-\epsilon,1+\epsilon\right)A^{\pi_{\theta_{k}}}(s,a)\right)\right],$ (5)
$A^{\pi_{\theta_{k}}}(s,a)=$ $A_{r}^{\pi_{\theta_{k}}}(s,a)\cdot I(s,s^{\prime})+A_{h}^{*}(s^{\textnormal{AV}},a^{\textnormal{AV}})\cdot(1-I(s,s^{\prime})),$ (6)
where $r_{t}(\theta)=\frac{\pi_{\theta}(a\mid s)}{\pi_{\theta_{k}}(a\mid s)}$ is the ratio, $A_{r}^{\pi_{\theta_{k}}}(s,a)=Q_{r}^{\pi_{\theta_{k}}}(s,a)-V_{r}^{\pi_{\theta_{k}}}(s)$ is CBV’s reward advantage, $A_{h}^{*}(s^{\textnormal{AV}},a^{\textnormal{AV}})$ is AV’s feasibility advantage, and $I(s,s^{\prime})=\mathbb{I}_{V^{\ast}_{h}(g(s))\leq 0,V_{h}^{\ast}(g(s^{\prime}))\leq 0}$ is the indicator.
According to Eq. 2, $Q_{h}^{*}(s^{\textnormal{AV}},a^{\textnormal{AV}})$ depends exclusively on $h(\cdot)$, which is in turn influenced by $s_{t}^{\textnormal{AV}},t\in\mathbb{N}$. The action $a^{\textnormal{AV}}$ affects only the next state ${s^{\textnormal{AV}}}^{\prime}$ through the environment transition function. Furthermore, when the BVs follow a deterministic policy, we can reasonably assume that the environment, excluding the CBV and AV, is deterministic. Therefore, we can express the next state of the AV as ${s^{\textnormal{AV}}}^{\prime}=P\left(s^{\textnormal{AV}},a^{\textnormal{AV}},s,a\right)$, where $P$ represents the deterministic environment transition function. This framework allows for the assertion that given the data sample, $\left(s^{\textnormal{AV}},a^{\textnormal{AV}},s,a\right)$, the next state of the AV is thereby determined, allowing us to derive the following lemma:
@@
As the BVs follow deterministic policy, the optimal feasible action-value function of AV can be achieved by AV’s current state and next state, while the action of the CBV and AV mainly influence state transition (see Section A.2 for proof).
$Q_{h}^{*}\left(s^{\textnormal{AV}},a^{\textnormal{AV}}\right)=\begin{cases}V_{h}^{*}(s^{\textnormal{AV}})&h({s^{\textnormal{AV}}}^{\prime})\geq h(s^{\textnormal{AV}})\\ \max\{h(s^{\textnormal{AV}}),V_{h}^{*}({s^{\textnormal{AV}}}^{\prime})\}&h({s^{\textnormal{AV}}}^{\prime})<h(s^{\textnormal{AV}})\end{cases}$ (7)
With Lemma 1 and the viewpoint transformation function $s^{\textnormal{AV}}=g\left(s\right)$, we can derive AV’s feasibility advantage through CBV’s current state $s$ and next state $s^{\prime}$ as:
$A_{h}^{*}(s^{\textnormal{AV}},a^{\textnormal{AV}})$ $=Q_{h}^{*}\left(s^{\textnormal{AV}},a^{\textnormal{AV}}\right)-V_{h}^{*}\left(s^{\textnormal{AV}}\right)$ (8)
=$=\begin{cases}V_{h}^{*}\left(g\left(s^{\prime}\right)\right)-V_{h}^{*}\left(g\left(s\right)\right)&h\left(g\left(s^{\prime}\right)\right)\geq h\left(g\left(s\right)\right)\\ \max\{h\left(g\left(s\right)\right),V_{h}^{*}\left(g\left(s^{\prime}\right)\right)\}-V_{h}^{*}\left(g\left(s\right)\right)&h\left(g\left(s^{\prime}\right)\right)<h\left(g\left(s\right)\right)\end{cases}$ (9)
By integrating Eqs. 5, 6 and 9, we formulate the online part of FREA in Algorithm 1.
@@
8:Online Part (Section 3.2)
9:Initialize policy parameters $\theta_{0}$, reward value function parameters $\psi_{0}$
@@
11:     Collect set of trajectories $\mathcal{B}_{k}=\{\tau_{i}\}$ with policy $\pi_{\theta_{k}}$, where $\tau_{i}$ is a $T$-step episode.